how many water molecules are there in ferrous sulphate crystal


1
Expert's answer
2021-06-21T05:31:02-0400

When water is used in the formation of crystals knows as water of crystallization. 'OR' 

Water of crystallization is the total weight of water in a substance at a given temperature and is mostly present in a definite ratio. 

On heating, iron(II) sulfate first loses its water of crystallization and the original green crystals are converted into a brown-colored anhydrous solid. 

FeSO4​.7H2

​O the heptahydrate in solution (water as solvent) transforms to both heptahydrate and tetrahydrate when the temperature reaches 56.6∘C.

= 7 × 6.023 ×1023 = 4.2161×1024 molecules